# The cohomological Kudla conjecture for unitary Shimura varieties

### v1

@@ 2.2. Homogeneous polynomials and modularity
    For $\ell,k<g+1$, we have $\Delta_{g+1}(m_{\ell,k})=m_{\ell,k}(P_{g})\leavevmode\nobreak\ ,$ by commutation of derivatives.
@@
    For $\ell,k<g+1$, using eq. 2.2, one can check that:
@@
    ${\frac{{}^{t}\partial(h(\lambda_{k},\lambda_{\ell}))}{\partial\overline{\lambda}_{g+1}}}\cdot\frac{\partial(m_{\ell,k}(P))}{\partial\lambda_{g+1}}=\delta_{g+1,\ell}\,\lambda_{k}\cdot\frac{\partial(m_{g+1,k}(P))}{\partial\lambda_{g+1}}=-P\leavevmode\nobreak\ ,$
@@
    If either $\ell=g+1$ or $k=g+1$ in the expressions in (4), then we get 00, as $m_{g+1,g+1}(P)$ does not depend neither on $\lambda_{g+1}$, nor on $\overline{\lambda_{g+1}}$.
@@
Finally, by subtracting eq. 2.5 from eq. 2.4, we get
$[\Lambda,\Delta]P=(2g-n)P=H(P)\leavevmode\nobreak\ ,$
@@
Let ${\mathcal{F}}_{n,g}^{\mathrm{prim}}=\mathrm{Ker}(\Delta_{g}:{\mathcal{F}}_{n,g}\rightarrow{\mathcal{F}}_{n,g-1})$. By virtue of Lemma 2.6, every polynomial in ${\mathcal{F}}_{n,g}$ is harmonic with respect all the variables, i.e., plurihamornic. By Theorem 2.9, we obtain a decomposition:
${\mathcal{F}}_{n,g}=\bigoplus_{\ell\leq g}\Lambda^{\ell}{\mathcal{F}}_{n,\ell}^{\mathrm{prim}}\leavevmode\nobreak\ .$
@@ 2.3. Projectors for $\mathfrak{sl}_{2}$ representations
By the results of the previous section, ${\mathcal{F}}_{n,\bullet}$ admits an action of the complex Lie algebra $\mathfrak{sl}_{2}$, with weights lying in the range $[-n,n]$.
In order to produce a correction to the generating series of compactified special cycles that is modular, we will use projectors onto the isotypic components of this $\mathfrak{sl}_{2}$ representation. This is normally done using Casimir elements, but for our application, we need only project vectors that are already in a fixed eigenspace for $H$, with eigenvalue $2g-n$, determined by the codimension of the special cycles. This simplifies the projector formulas that we recall in the general setting.
@@
where $\pi_{k}\simeq\mathrm{Sym}^{k}(V_{std})$ is the unique irreducible representation of dimension $k+1$. Here, $V_{std}$ is the defining representation of $\mathfrak{sl}_{2}$, and $U_{k}$ is a trivial representation that encodes the multiplicity with which $\pi_{k}$ occurs in $W$. The eigenspaces of $H$ acting on $\pi_{k}$ are each one-dimensional, with integral eigenvalues called weights in $[-i,i]$. Any two weights for $\pi_{k}$ differ by an even integer.
@@
Furthermore, $\rho(\Pi_{m,k})$ is an isotypic projector when restricted to $(-m)$-eigenspaces.
@@
For a $(-m)$-eigenvector $v\in W$ as above, its isotypic decomposition only involves nonzero summands in $\pi_{k}\otimes U_{k}$ for $k\geq m$ and $k\equiv m$ (mod 2), so we write:
@@
where $v_{k}\in\pi_{k}\otimes U_{k}$ has weight $-m$. The first summand $v_{m}$ will coincide with the primitive part of $v$. By a computation in the irreducible representation $\pi_{k}$, we have
@@ 2.4. Theta series and Hermitian (quasi)-modular forms
We recall now a useful modularity criterion for Hermitian modular forms due to Shimura [Shi97, Appendix 7]. We keep all notation from the previous section.
Let $k$ be an imaginary quadratic field and let $(L,h)$ be a $\mathcal{O}_{k}$-Hermitian module of signature $(n,0)$, with associated Weil representation $\rho_{L,g}$ for $g\geq 1$.
We have an identification $(L_{\mathbb{R}},h)\simeq(\mathbb{C}^{n},h_{st})$ where $h_{st}$ is the standard Hermitian form on $\mathbb{C}^{n}$. This identification is obtained as follows: let $H$ be the matrix of the Hermitian pairing $h$ in a $k$-basis of $L$ and let $\sqrt{H}$ be the unique Hermitian positive definite matrix such that $\sqrt{H}^{2}=H$: then the above identification is given by $\lambda\mapsto\sqrt{{}^{t}H}^{-1}\lambda$ Therefore, the differential operators considered in Section 2.2 are written in the coordinates of $(L_{\mathbb{R}},h)$ as follows:
$\Delta_{i}=\prescript{t}{}{\left(\frac{\partial}{\partial\lambda_{i}}\right)}\cdot H^{-1}\cdot\frac{\partial}{\partial\overline{\lambda_{i}}}\leavevmode\nobreak\ ,$
for $\lambda_{i}\in L_{\mathbb{R}}$, and similarly for the directional derivative, which becomes ${}^{t}\lambda\cdot H^{-1}\cdot\frac{\partial}{\partial\mu}$, for $\lambda,\mu\in L_{\mathbb{R}}$ written with respect to a basis of $L$.
@@
$\vartheta_{P}(\tau)=\det(Y)^{-1}\sum_{\underline{\lambda}\in(L^{\vee})^{g}}\exp\left(-\frac{\Delta}{4\pi}\right)\left(P\right)(\underline{\lambda}\cdot Y^{\frac{1}{2}})q^{h(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\in\mathbb{C}[(L^{\vee}/L)^{g}]\llbracket q\rrbracket\leavevmode\nobreak\ ,$
transforms like a Hermitian modular form of weight $2+n$ in the variable $\tau=X+iY\in\mathcal{H}_{g}$, with respect to the Weil representation $\rho_{L,g}$ of $\mathrm{U}(g,g)(\mathbb{Z})$
@@
$\vartheta_{P}^{+}=\sum_{\underline{\lambda}\in(L^{\vee}/L)^{g}}P(\underline{\lambda})q^{Q(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\leavevmode\nobreak\ .$
Therefore, the theorem provides natural non-holomorphic completions that transform like Hermitian modular forms. We then make the follow definition.
@@
The holomorphic function $\vartheta_{P}^{+}$ is called a Hermitian quasi-modular form. The span of all such functions will be denoted $\mathrm{QHerMod}(2+n,\rho_{L,g})$.
2.5. Correction of theta series
@@
$0=X\cdot(\psi\otimes f)=-(\psi\circ X)\otimes f+\psi\otimes(X\cdot f)\leavevmode\nobreak\ .$
@@
Let $2g\leq n$, and let $m=n-2g$. Since $u$ is a morphism of $\mathfrak{sl}_{2}$-representations, we get a morphism
$u_{g}:V_{-m}\rightarrow{\mathcal{F}}_{n,g}\leavevmode\nobreak\ ,$
@@
$\left[\Pi_{m,k}^{*}\otimes 1\right](u_{g})=\left[1\otimes\Pi_{m,k}\right](u_{g})\in(V^{*})_{\pi_{k},m}\otimes({\mathcal{F}}_{n,g})_{\pi_{k}}\leavevmode\nobreak\ ,$
@@
$u_{g}=\sum_{k\geq m}[\Pi^{*}_{m,k}\otimes 1](u_{g})=\sum_{k\geq m}[1\otimes\Pi_{m,k}](u_{g})\leavevmode\nobreak\ ,$
@@
$\Phi_{g}=\sum_{\underline{\lambda}\in(L^{\vee})^{g}}u_{g}(\lambda_{1},\ldots,\lambda_{g})q^{h(\underline{\lambda})}{[\underline{\lambda}]}\in(V^{*})_{m}\otimes\mathbb{C}[(L^{\vee}/L)^{g}]\llbracket q\rrbracket\leavevmode\nobreak\ ,$
which, by Theorem 2.12, transforms like a Hermitian quasi-modular form. Write the Lefschetz decompositions:
$V^{*}_{m}=\bigoplus_{k=m+2r}(E^{*})^{r}V^{*}_{\pi_{k},k}\quad\textrm{and}\quad{\mathcal{F}}_{n,g}=\bigoplus_{0\leq r\leq g}\Lambda^{r}{\mathcal{F}}_{n,g-r}^{\mathrm{prim}}\leavevmode\nobreak\ .$
@@
$[\Pi_{m,m}^{*}\otimes\vartheta](u_{g})=\Phi_{g}-\sum_{{k>m}}[\Pi_{m,k}^{*}\otimes\vartheta](u_{g})$
@@
Indeed, $[\Pi_{m,m}^{*}\otimes\vartheta](u_{g})\in V^{*}_{\pi_{m},m}\otimes{\mathcal{F}}_{n,g}^{\mathrm{prim}}$ therefore applying the theta lifts produces holomorphic modular forms of weight $2+n$ with respect to the Weil representation $\rho_{L,g}$ of $\mathrm{U}(g,g)(\mathbb{Z})$.
@@
Observe that for $k>m$, $[\Pi_{m,k}^{*}\otimes\vartheta](u_{g})$ is in the image of $E^{*}\otimes 1$ in $V^{*}\otimes\mathrm{QHerMod}(2+n,\rho_{L,g})$, which will be useful for corrections by the boundary components later, as $E^{*}$ will be the cup product with (a multiple of) the Chern class of the conormal bundle to the boundary.
@@ 3.1. Unitary Shimura varieties
${\mathbf{G}}(R)=\{g\in\mathrm{End}_{k}(V)\otimes_{\mathbb{Q}}R\,|\,gg^{*}=\mu(g)\in R^{\times}\}\leavevmode\nobreak\ ,$
@@
${\mathbb{D}}(V)\simeq\mathrm{U}(n+1,1)/\mathrm{U}(n+1)\times\mathrm{U}(1)\leavevmode\nobreak\ ,$
@@ 3.2. Special cycles and modularity
Let $11\leq g\leq n+$ 1. For $g$ vectors $\underline{\lambda}=(\lambda_{1},\ldots,\lambda_{g})$, we denote their Gram matrix by $h(\underline{\lambda})$. Let $\mathbb{D}_{\underline{\lambda}}(V)\subset\mathbb{D}(V)$ be the closed complex analytic subspace defined by the following conditions on $P\in\mathbb{D}(V)$: $\lambda_{i}\in P^{\bot}$ for all $1\leq i\leq g$. Since $P^{\bot}$ is a positive definite Hermitian lattice, $\mathbb{D}_{\underline{\lambda}}(V)$ is empty unless $h(\underline{\lambda})$ is a positive semidefinite Hermitian matrix. In that case, $\mathbb{D}_{\underline{\lambda}}(V)$ is a Hermitian symmetric subdomain of codimension equal to the rank of the matrix $h(\underline{\lambda})$. Let $N\in\mathrm{Herm}_{g}(k)$ be a $g\times g$ Hermitian positive semidefinite matrix. We define the following cycle:
${\mathcal{Z}}(\underline{\nu},N)=\Gamma_{L}\big{\backslash}\left(\bigcup_{\underset{h(\underline{\lambda})=N}{\underline{\lambda}\in\underline{\nu}+L^{g}}}\mathbb{D}(V)_{\underline{\lambda}}\right)\leavevmode\nobreak\ .$
Then ${\mathcal{Z}}(\underline{\nu},N)\hookrightarrow X_{\Gamma}$ is a closed algebraic subvariety of codimension $r(N)$, the rank of the matrix $N$. Let $\mathbb{E}$ be the tautological line bundle on $X_{\Gamma}$ and define the generating series:
(3.1) $\Phi_{L}(\tau)=\sum_{\underset{N\in\mathrm{Herm}_{g}(k)_{\geq 0}}{\underline{\nu}\in(L^{\vee}/L)^{g}}}[{\mathcal{Z}}(\underline{\nu},N)]\cup c_{1}(\mathbb{E}^{\vee})^{g-r(N)}q^{N}\mathfrak{e}_{\underline{\nu}}\in\mathrm{CH}^{g}(X_{\Gamma})\otimes\mathbb{C}[(L^{\vee}/L)^{g}]\llbracket q\rrbracket,$

### v2

@@ 2.3. Homogeneous polynomials and modularity
    For $\ell,k<g+1$, we have $\Delta_{g+1}(m_{\ell,k})=m_{\ell,k}(P_{g})\penalty 10000\ ,$ by commutation of derivatives.
@@
    For $\ell,k<g+1$, using eq.˜2.2, one can check that:
@@
    ${\frac{{}^{t}\partial(h(\lambda_{k},\lambda_{\ell}))}{\partial\overline{\lambda}_{g+1}}}\cdot\frac{\partial(m_{\ell,k}(P))}{\partial\lambda_{g+1}}=\delta_{g+1,\ell}\,\lambda_{k}\cdot\frac{\partial(m_{g+1,k}(P))}{\partial\lambda_{g+1}}=-P\penalty 10000\ ,$
@@
    If either $\ell=g+1$ or $k=g+1$ in the expressions in (4), then we get 0, as $m_{g+1,g+1}(P)$ does not depend neither on $\lambda_{g+1}$, nor on $\overline{\lambda_{g+1}}$.
@@
Finally, by subtracting eq.˜2.5 from eq.˜2.4, we get
$\Lambda,\Delta]P=(2g-n)P=H(P)\penalty 10000\ ,$
@@
Let ${\mathcal{F}}_{n,g}^{\mathrm{prim}}=\mathrm{Ker}(\Delta_{g}:{\mathcal{F}}_{n,g}\rightarrow{\mathcal{F}}_{n,g-1})$. By virtue of Lemma˜2.6, every polynomial in ${\mathcal{F}}_{n,g}^{\mathrm{prim}}$ is harmonic with respect to all the variables, i.e., pluriharmonic. By Theorem˜2.9, we obtain a decomposition:
${\mathcal{F}}_{n,g}=\bigoplus_{\ell\leq g}\Lambda^{\ell}{\mathcal{F}}_{n,\ell}^{\mathrm{prim}}\penalty 10000\ .$
@@ 2.4. Projectors for $\mathfrak{sl}_{2}$ representations
By the results of the previous section, ${\mathcal{F}}_{n,\bullet}$ admits an action of the complex Lie algebra $\mathfrak{sl}_{2}$, with weights lying in the range [n,n][-n,n].
In order to produce a correction to the generating series of compactified special cycles that is modular, we will use projectors onto the isotypic components of this $\mathfrak{sl}_{2}$ representation. This is normally done using Casimir elements, but for our application, we need only project vectors that are already in a fixed eigenspace for $H$, with eigenvalue 2gn2g-n, determined by the codimension of the special cycles. This simplifies the projector formulas that we recall in the general setting.
@@
where $\pi_{k}\simeq\mathrm{Sym}^{k}(V_{std})$ is the unique irreducible representation of dimension $k+1$. Here, $V_{std}$ is the defining representation of $\mathfrak{sl}_{2}$, and $U_{k}$ is a trivial representation that encodes the multiplicity with which $\pi_{k}$ occurs in $W$. The eigenspaces of $H$ acting on $\pi_{k}$ are each one-dimensional, with integral eigenvalues called weights in [i,i][-i,i]. Any two weights for $\pi_{k}$ differ by an even integer.
@@
Furthermore, $\rho(\Pi_{m,k})$ is an isotypic projector when restricted to (m)(-m)-eigenspaces.
@@
For a (m)(-m)-eigenvector $v\in W$ as above, its isotypic decomposition only involves nonzero summands in $\pi_{k}\otimes U_{k}$ for $k\geq m$ and $k\equiv m$ (mod 2), so we write:
@@
where $v_{k}\in\pi_{k}\otimes U_{k}$ has weight m-m. The first summand $v_{m}$ will coincide with the primitive part of $v$. By a computation in the irreducible representation $\pi_{k}$, we have
@@ 2.5. Theta series and Hermitian (quasi)-modular forms
We now introduce the notion of Hermitian quasi-modularity and a useful criterion in the case of weighted theta functions. We keep all notation from the previous section. Let $k$ be an imaginary quadratic field, and let $(M,h)$ be an $\mathcal{O}_{k}$-module with a Hermitian form of signature $(n,0)$ and the associated Weil representation $\rho_{M,g}$ for $g\geq 1$.
We have an identification $(M_{\mathbb{R}},h)\simeq(\mathbb{C}^{n},h_{st})$ where $h_{st}$ is the standard Hermitian form on $\mathbb{C}^{n}$. This identification is obtained as follows: let $H$ be the matrix of the Hermitian pairing $h$ in a $k$-basis of $M$ and let $\sqrt{H}$ be the unique Hermitian positive definite matrix such that $\sqrt{H}^{2}=H$: then the above identification is given by $\lambda\mapsto\sqrt{{}^{t}H}^{-1}\lambda$ . Therefore, the differential operators considered in Section˜2.3 are written in the coordinates of $(M_{\mathbb{R}},h)$ as follows:
$\Delta_{i}=\prescript{t}{}{\left(\frac{\partial}{\partial\lambda_{i}}\right)}\cdot H^{-1}\cdot\frac{\partial}{\partial\overline{\lambda_{i}}}\penalty 10000\ ,$
for $\lambda_{i}\in M_{\mathbb{R}}$, and similarly for the directional derivative, which becomes $^{t}\lambda\cdot H^{-1}\cdot\frac{\partial}{\partial\mu}$, for $\lambda,\mu\in M_{\mathbb{R}}$ written with respect to a basis of $M$.
The following theorem is the Hermitian analogue of a theorem of Roehrig [Roe21] in the orthogonal case. A detailed proof will appear in the upcoming work of Ben Howard.
@@
$\vartheta_{P}(\tau)=\det(Y)^{-1}\sum_{\underline{\lambda}\in(M^{\vee})^{g}}\exp\left(-\frac{\Delta}{4\pi}\right)\left(P\right)(\underline{\lambda}\cdot Y^{\frac{1}{2}})q^{h(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\in\mathbb{C}[(M^{\vee}/M)^{g}]\llbracket q\rrbracket\penalty 10000\ ,$
transforms like a Hermitian modular form of weight $2+n$ in the variable $\tau=X+iY\in\mathcal{H}_{g}$, with respect to the Weil representation $\rho_{M,g}$ of $\mathrm{U}(g,g)(\mathbb{Z})$ .
@@
$\vartheta_{P}^{+}=\sum_{\underline{\lambda}\in(M^{\vee})^{g}}P(\underline{\lambda})q^{Q(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\penalty 10000\ .$
Therefore, the theorem provides natural non-holomorphic completions that transform like Hermitian modular forms. We make the following more general definition.
@@
Let $f:{\mathcal{H}}_{g}\rightarrow\mathbb{C}[(M^{\vee}/M)^{g}]$ be a holomorphic function. We say that $f$ is a Hermitian holomorphic quasi-modular form of weight $k$ with respect to $\rho_{M,g}$ if there exist holomorphic functions $g_{1},\ldots,g_{r}$ and polynomials $P_{j}$ in $\det(Y)^{-1}$ and the entries of $Y^{\frac{1}{2}}$, where $Y=\frac{\tau^{*}-\tau}{2i}$, such that
$\widehat{f}=f+\sum_{j=1}^{r}g_{j}P_{j}$
transforms like a modular form of weight $k$ with respect to $\rho_{M,g}$
Example 2.14.
The holomorphic function $\vartheta_{P}^{+}$ is an example of a Hermitian quasi-modular form. The span of all such functions will be denoted
$\mathrm{QHerMod}(2+n,\rho_{M,g}).$
2.6. Corrections of theta series
@@
$0=X\cdot(\psi\otimes f)=-(\psi\circ X)\otimes f+\psi\otimes(X\cdot f)\penalty 10000\ .$
@@
Let $2g\leq n$, and let m=n2gm=n-2g. Since $u$ is a morphism of $\mathfrak{sl}_{2}$-representations, we get a morphism
$u_{g}:V_{-m}\rightarrow{\mathcal{F}}_{n,g}\penalty 10000\ ,$
@@
$\left[\Pi_{m,k}^{*}\otimes 1\right](u_{g})=\left[1\otimes\Pi_{m,k}\right](u_{g})\in(V^{*})_{\pi_{k},m}\otimes({\mathcal{F}}_{n,g})_{\pi_{k}}\penalty 10000\ ,$
@@
$u_{g}=\sum_{k\geq m}[\Pi^{*}_{m,k}\otimes 1](u_{g})=\sum_{k\geq m}[1\otimes\Pi_{m,k}](u_{g})\penalty 10000\ ,$
@@
$\Phi_{g}=\sum_{\underline{\lambda}\in(L^{\vee})^{g}}u_{g}(\lambda_{1},\ldots,\lambda_{g})q^{h(\underline{\lambda})}{[\underline{\lambda}]}\in(V^{*})_{m}\otimes\mathbb{C}[(L^{\vee}/L)^{g}]\llbracket q\rrbracket\penalty 10000\ ,$
which, by Theorem˜2.12, transforms like a Hermitian quasi-modular form. Write the Lefschetz decompositions:
$V^{*}_{m}=\bigoplus_{k=m+2r}(E^{*})^{r}V^{*}_{\pi_{k},k}\quad\textrm{and}\quad{\mathcal{F}}_{n,g}=\bigoplus_{0\leq r\leq g}\Lambda^{r}{\mathcal{F}}_{n,g-r}^{\mathrm{prim}}\penalty 10000\ .$
@@
$\Pi_{m,m}^{*}\otimes\vartheta](u_{g})=\Phi_{g}-\sum_{{k>m}}[\Pi_{m,k}^{*}\otimes\vartheta](u_{g})$
@@
Indeed, $\Pi_{m,m}^{*}\otimes\vartheta](u_{g})\in V^{*}_{\pi_{m},m}\otimes{\mathcal{F}}_{n,g}^{\mathrm{prim}}$ therefore applying the theta lifts produces holomorphic modular forms of weight $2+n$ with respect to the Weil representation $\rho_{L,g}$ of $\mathrm{U}(g,g)(\mathbb{Z})$.
@@
Observe that for $k>m$, $\Pi_{m,k}^{*}\otimes\vartheta](u_{g})$ is in the image of $E^{*}\otimes 1$ in $V^{*}\otimes\mathrm{QHerMod}(2+n,\rho_{L,g})$, which will be useful for corrections by the boundary components later, as $E^{*}$ will be the cup product with (a multiple of) the Chern class of the conormal bundle to the boundary.
@@ 3.1. Unitary Shimura varieties
${\mathbf{G}}(R)=\{g\in\mathrm{End}_{k}(V)\otimes_{\mathbb{Q}}R\,|\,gg^{*}=\mu(g)\in R^{\times}\}\penalty 10000\ ,$
@@
${\mathbb{D}}(V)\simeq\mathrm{U}(n+1,1)/\mathrm{U}(n+1)\times\mathrm{U}(1)\penalty 10000\ ,$
@@ 3.2. Special cycles and modularity
Let $1\leq g\leq n+1$. For $g$ vectors $\underline{\lambda}=(\lambda_{1},\ldots,\lambda_{g})$, we denote their Gram matrix by $h(\underline{\lambda})$. Let $\mathbb{D}_{\underline{\lambda}}(V)\subset\mathbb{D}(V)$ be the closed complex analytic subspace defined by the following conditions on $P\in\mathbb{D}(V)$: $\lambda_{i}\in P^{\bot}$ for all $1\leq i\leq g$. Since $P^{\bot}$ is a positive definite Hermitian lattice, $\mathbb{D}_{\underline{\lambda}}(V)$ is empty unless $h(\underline{\lambda})$ is a positive semidefinite Hermitian matrix. In that case, $\mathbb{D}_{\underline{\lambda}}(V)$ is a Hermitian symmetric subdomain of codimension equal to the rank of the matrix $h(\underline{\lambda})$. Let $N\in\mathrm{Herm}_{g}(k)$ be a $g\times g$ Hermitian positive semidefinite matrix. We define the following cycle:
${\mathcal{Z}}^{\rm naive}(\underline{\nu},N)=\Gamma_{L}\big\backslash\left(\bigcup_{\underset{h(\underline{\lambda})=N}{\underline{\lambda}\in\underline{\nu}+L^{g}}}\mathbb{D}(V)_{\underline{\lambda}}\right)\penalty 10000\ .$
Then ${\mathcal{Z}}^{\rm naive}(\underline{\nu},N)\hookrightarrow X_{\Gamma}$ is a closed algebraic subvariety of codimension $r(N)$, the rank of the matrix $N$. Let $\mathbb{E}$ be the tautological line bundle on $X_{\Gamma}$. To ensure that all the special cycles live in the same codimension, we define:
${\mathcal{Z}}(\underline{\nu},N)={\mathcal{Z}}^{\rm naive}(\underline{\nu},N)\cup c_{1}(\mathbb{E}^{\vee})^{g-r(N)}\penalty 10000\ .$
Define the generating series [^3] It is an open problem whether this series is convergent with values in the Chow group.:
(3.1) $\Phi^{g}_{L}(\tau)=\sum_{\underset{N\in\mathrm{Herm}_{g}(k)_{\geq 0}}{\underline{\nu}\in(L^{\vee}/L)^{g}}}[{\mathcal{Z}}(\underline{\nu},N)]q^{N}\mathfrak{e}_{\underline{\nu}}\in\mathrm{CH}^{g}(X_{\Gamma})\otimes\mathbb{C}[(L^{\vee}/L)^{g}]\llbracket q\rrbracket,$